# DiffPF: Differentiable Particle Filtering with Generative Sampling via Conditional Diffusion Models

### v1

@@ Abstract
This paper proposes DiffPF, a differentiable particle filter that leverages diffusion models for state estimation in dynamic systems. Unlike conventional differentiable particle filters—which require importance weighting and typically rely on predefined or low-capacity proposal distributions. DiffPF learns a flexible posterior sampler by conditioning a diffusion model on predicted particles and the current observation. This enables accurate, equally-weighted sampling from complex, high-dimensional, and multimodal filtering distributions. We evaluate DiffPF across a range of scenarios, including both unimodal and highly multimodal distributions, and test it on simulated as well as real-world tasks, where it consistently outperforms existing filtering baselines. In particular, DiffPF achieves an 82.8% improvement in estimation accuracy on a highly multimodal global localization benchmark, and a 26% improvement on the real-world KITTI visual odometry benchmark, compared to state-of-the-art differentiable filters. To the best of our knowledge, DiffPF is the first method to integrate conditional diffusion models into particle filtering, enabling high-quality posterior sampling that produces more informative particles and significantly improves state estimation. The code will be made available online upon acceptance.
@@ I INTRODUCTION
Estimating the hidden state of a dynamic system from partial and noisy observations is a core challenge across a wide range of applications, including navigation [1], control [2], and robotics [3]. Kalman filters and their extensions [4] have long served as standard tools within the Bayesian filtering framework. However, their reliance on linearity and Gaussian noise assumptions severely limits their applicability to real-world systems that are often nonlinear, high-dimensional and exhibit multimodal behaviors.
@@
An alternative line of work, particle filters, avoids these restrictive assumptions by representing the posterior distribution with a set of weighted samples instead of relying on closed-form expressions. Bootstrap particle filter [5], one of the most commonly used algorithms, draws particles from the transition model and assigns weights based on observation likelihoods. Rao-Blackwellized particle filter [6] further improves efficiency by analytically marginalizing part of the latent state using parametric inference, thereby reducing sampling variance. However, these methods still rely on manually designed or limited-capacity proposals and are prone to sample degeneracy, particularly in high-dimensional or perceptually complex domains.
@@
Recent developments in deep learning have introduced data-driven alternatives to traditional state estimation methods. Deep state-space models (DSSMs) [7] provide a unified framework capable of capturing nonlinear latent dynamics and processing high-dimensional inputs, making them particularly suited for complex real-world environments. Differentiable Particle Filters (DPFs), a class of methods that embed sample-based filtering into DSSMs [8, 9], enable end-to-end learning of recursive Bayesian inference over learned dynamics and observation models. However, most DPFs still inherit several limitations from classical particle filtering. The proposal distributions used in DPFs are often either hand-crafted or limited in capacity, making it difficult to match the true filtering posterior [10]. Furthermore, they rely on importance sampling, which necessitates explicit weighting of particles—a process that becomes increasingly unreliable in high-dimensional or highly multimodal settings due to particle degeneracy.
@@
To overcome these challenges, we employ diffusion models as powerful implicit samplers, capable of capturing rich and multimodal posterior distributions. This allows for a more expressive and learnable formulation of the state update step within the filtering process. Specifically, DiffPF employs a diffusion model for state update, conditioned on both the predicted particles and the current observation to generate refined posterior samples. An example application of DiffPF on the KITTI visual odometry benchmark [11] is shown in Figure LABEL:fig:teaser. DiffPF departs from particle filtering by eliminating the reliance on importance sampling and explicit proposal distributions. Instead, it leverages a diffusion model to sample directly from the filtering posterior, enabling accurate inference without the need for particle weighting. This sampling-based yet weight-free approach naturally avoids particle degeneracy and, consequently, eliminates the need for the non-differentiable resampling step, while supporting multimodal and high-dimensional distributions.
The main contributions of this paper are summarized as follows:
@@ II-A Particle filter
Particle filters (PFs) are a class of sequential Monte Carlo (SMC) methods widely used for recursive state estimation in nonlinear and non-Gaussian systems [12]. Classical particle filters, such as the Bootstrap Filter[5], are built on the Sampling-Importance-Resampling (SIR) framework [13], which approximates the filtering distribution using weighted particles. While theoretically general, these methods often struggle in practice due to particle degeneracy—where most weights collapse to near zero—and the difficulty of designing effective proposal distributions, especially in high-dimensional or perceptually complex environments.
Several extensions have been proposed to improve the sampling efficiency of particle filters. General importance sampling filters [14] aim to use more informed proposal distributions that better align with the true posterior, while the Auxiliary Particle Filter (APF) [15] incorporates observation information to guide the selection of particles before propagation. In parallel, sequential MCMC methods [16] replace the standard sampling step in particle filters with Markov chain updates over full or partial trajectories, allowing for posterior-consistent sampling. Particle MCMC (PMCMC)[17] integrates particle filtering into Markov chain Monte Carlo algorithms for joint parameter and state inference. Detailed treatments of these and other developments in particle filtering can be found in [13].
@@ II-B Differentiable Particle Filters
The aforementioned methods often assume known system dynamics or observation models, which often do not hold in real-world scenarios [18]. Differentiable particle filters [8, 9, 10, 19, 20, 21, 22, 23] address this by integrating neural networks into the filtering process, enabling end-to-end learning from data.
Most DPFs model transition dynamics by mapping previous states to the next latent state using either fully connected networks[9, 10], recurrent architectures such as LSTMs and GRUs[20, 23], or fixed parametric forms with learnable noise parameters[22]. Observation models are typically implemented as neural networks that either output the parameters of a predefined distribution (e.g., the mean and variance of a Gaussian) [23] or produce unnormalized scores interpreted as pseudo-likelihoods [19]. While effective in simple scenarios, these designs are often restricted to specific distribution families and struggle to represent complex or multimodal posteriors, limiting the expressiveness of the resulting filtering distributions and motivating the need for a more flexible and general framework.
The proposal distribution, which guides particle sampling, is typically handcrafted [24, 25] or implemented as a neural network conditioned on the current observation and possibly past states. In many DPFs, the transition model is reused as the proposal, and particles are often generated deterministically via regression rather than sampling, which leads to particle degeneracy [26], reduces particle quality, and constrains the representational capacity of the filtering distribution. Although [18] proposes using normalizing flows to sample from the proposal distribution, the method remains confined to the importance sampling framework and suffers from accumulated estimation errors due to sampling from multiple distributions simultaneously. To enable gradient-based learning, DPFs must address the non-differentiability of traditional resampling operations. Several methods replace hard resampling with differentiable approximations [27], such as soft-resampling [10] and entropy-regularized optimal transport solvers [19]. While these approximations allow gradients to flow through resampling steps, they often introduce additional computational complexity or require task-specific tuning, and particle degeneracy can still persist.
Figure 0: At time $t$, the estimated particles from time $t{-}1$ are propagated through the process model to obtain a prior distribution over the current state. Simultaneously, the observation $\bm{o}_{t}$ is encoded into a feature representation. These two sources of information are jointly used as the condition for a conditional diffusion model, which iteratively refines a set of noisy latent samples to generate equally-weighted particles that approximate the filtering posterior.
@@ II-C Diffusion Models for Probabilistic Modeling
Recent advances in generative modeling have highlighted diffusion models as effective tools for capturing intricate, multimodal distributions across a range of data types. These models operate by progressively denoising latent variables [28, 29], making them particularly adept at representing uncertainty and complex data structures.
In the domain of robotics, such approaches as Diffusion Policy [30] and NoMaD [31] have demonstrated the effectiveness of conditional diffusion models in generating action sequences by capturing the inherently multi-modal nature of action distributions [32]. Meanwhile, diffusion models have gained significant attention for their effectiveness in trajectory generation [33, 34]. Building on the success of these tasks, we explore their potential in state estimation. Existing attempts include introducing diffusion models into Kalman filters [35]. We further develop a novel sample-based framework that integrates diffusion models into differentiable particle filtering, and propose end-to-end efficient training algorithms to facilitate the filter design. To our knowledge, this is the first such integration, and it significantly improves estimation accuracy across a variety of scenarios.
@@ III Method
In sequential inference, a Bayesian filter aims to recursively estimate the current state $\bm{x}_{t}$ via the posterior distribution $\text{post}(\bm{x}_{t}):=p(\bm{x}_{t}\mid\bm{o}_{1:t},\bm{a}_{1:t})$, conditioned on all observations and actions up to time $t$. Under the Markov assumption, all relevant information from past observations $\bm{o}_{1:t}$ and actions $\bm{a}_{1:t}$ is captured by the current posterior $\text{post}(\bm{x}_{t})$. By exploiting the conditional independence of observations and actions given the latent state, the inference proceeds recursively in two steps [36, 9]. In the prediction step, the posterior from the previous timestep is propagated through the system’s transition dynamics to yield a prior over the current state:
[…]
This predicted prior $p(\bm{x}_{t}\mid\bm{o}_{1:t-1},\bm{a}_{1:t})$ implicitly reflects all relevant past information, as conveyed by the previous posterior and the process model. In the update step, the filter updates its estimate of the current state by integrating the prior distribution, obtained from the prediction step, with the latest observation $\bm{o}_{t}$. As illustrated in Figure II-B, our method follows this two-step Bayesian filtering procedure and represents the posterior $\text{post}(\bm{x}_{t})$ using a set of equally weighted particles $\bm{\chi}_{t}:=\{\bm{x}_{t}^{(i)}\}_{i=1}^{N}$. The remainder of this section details each component of the proposed framework.
III-A System and Perception Modeling
Given a set of $N$ particles $\{\bm{x}_{t-1}^{(i)}\}_{i=1}^{N}$ representing the posterior at time $t{-}1$, we apply a process model to obtain predicted particles ${\hat{\bm{\chi}}_{t}}:=\{\hat{\bm{x}}_{t}^{(i)}\}_{i=1}^{N}$ for the current step. Specifically, each particle is propagated through a learnable or given process model $f_{\text{dyn}}$ as:
@@ III-A System and Perception Modeling
where $\bm{a}_{t}$ is the action at time $t$. Through this step, the relevant information from the action sequence $\bm{a}_{1:t}$ and past observations $\bm{o}_{1:t-1}$ is encoded into the predicted particle set $\{\hat{\bm{x}}_{t}^{(i)}\}_{i=1}^{N}$. The process model $f_{\text{dyn}}$ can be defined using either parametric forms based on system knowledge, or learned from data using neural networks such as Multi-Layer Perceptrons(MLPs), stochastic models [37], or generative approaches like normalizing flows [38].
@@ III-B Update Step
Update step refines the prior belief over the current state by incorporating the latest observation. In conventional DPFs, the update is typically implemented by assigning importance weights to particles based on an observation model, followed by resampling to mitigate particle degeneracy. However, these procedures often require carefully designed likelihood functions and involve non-differentiable operations. In contrast, our method performs the update implicitly through a conditional diffusion model. Specifically, we make a assumption that
@@
This assumption is justified because $\bm{c}_{t}$ captures all temporal information from $\bm{o}_{1:t}$ and $\bm{a}_{1:t}$. Instead of sampling directly from the intractable posterior distribution $p(\bm{x}_{t}\mid\bm{o}_{1:t},\bm{a}_{1:t})$, we choose to sample from the approximated distribution $p(\bm{x}_{t}\mid\bm{c}_{t})$. This yields a set of particles $\{\bm{x}_{t}^{(i)}\}_{i=1}^{N}$, from which we can derive:
$\text{post}(\bm{x}_{t})$ $=\eta\sum_{i=1}^{N}\frac{p(\bm{x}_{t}^{i}\mid\bm{o}_{1:t},\bm{a}_{1:t})}{p(\bm{x}_{t}^{i}\mid\bm{c}_{t})}\delta(\bm{x}_{t}-\bm{x}_{t}^{(i)}),$ (6)
$\approx\eta\sum_{i=1}^{N}\delta(\bm{x}_{t}-\bm{x}_{t}^{(i)})=\frac{1}{N}\sum_{i=1}^{N}\delta(\bm{x}_{t}-\bm{x}_{t}^{(i)}),$
where $\delta(\cdot)$ denotes the Dirac function and $\eta$ is a normalization constant. This indicates that all sampled particles share equal weights, thereby eliminating the need for weighting or resampling.
Diffusion models have been shown to approximate arbitrarily complex probability distributions by learning time-dependent score functions that guide the reversal of a noise-perturbed process[39]. In our case, a conditional diffusion model is trained to approximate $p(\bm{x}_{t}\mid\bm{c}_{t})$ and generate particle samples. Specifically, we instantiate a reverse diffusion process with a Denoising Diffusion Probabilistic Model (DDPM), which iteratively refines a random Gaussian sample toward the target posterior distribution. We begin by drawing an initial Gaussian sample:
@@
where $K$ is the total number of reverse diffusion steps. At each step $k$, the model predicts the noise component $\bm{\epsilon}_{\theta}(\bm{x}_{t,k}^{(i)},k,\bm{c}_{t})$, which subsequently serves to infer a cleaner approximation of the sample at timestep $k-1$:
@@
$:=\beta(\bm{x}_{t,k}^{(i)}+\gamma\epsilon_{\theta}(\bm{x}_{t,k}^{(i)},k,\bm{c}_{t}))+\sigma\bm{z}.$
Here $\bm{z}\sim\mathcal{N}(0,I)$, and $\bm{x}_{t,k}^{(i)}$ refers to the $i$-th corrupted sample at diffusion step $k$. $\alpha$, $\bar{\alpha}$, ${\sigma}$ are fixed noise schedule parameters as defined in [40], while $\beta$ and $\gamma$ are scaling terms introduced for simplicity. This stochastic refinement continues recursively until $k=0$, yielding the final particle sample
$\bm{x}_{t,0}^{(i)}=\bm{x}_{t}^{(i)}.$ (9)
This reverse diffusion process is repeated $N$ times to produce the final particle set $\bm{\chi}_{t}=\{\bm{x}_{t}^{(i)}\}_{i=1}^{N}$. As all particles are equally weighted, the final estimate is given by the average over the particle set
$\bm{x}_{t}=\frac{1}{N}\sum_{i=1}^{N}\bm{x}_{t}^{(i)}$ (10)

### v2

@@ Abstract
This paper proposes DiffPF, a differentiable particle filter that leverages diffusion models for state estimation in dynamic systems. Unlike conventional differentiable particle filters, which require importance weighting and typically rely on predefined or low-capacity proposal distributions, DiffPF learns a flexible posterior sampler by conditioning a diffusion model on predicted particles and the current observation. This enables accurate, equally-weighted sampling from complex, high-dimensional, and multimodal filtering distributions. We evaluate DiffPF across a range of scenarios, including both unimodal and highly multimodal distributions, and test it on simulated as well as real-world tasks, where it consistently outperforms existing filtering baselines. In particular, DiffPF achieves a 90.3% improvement in estimation accuracy on a highly multimodal global localization benchmark, and a nearly 50% improvement on the real-world robotic manipulation benchmark, compared to state-of-the-art differentiable filters. To the best of our knowledge, DiffPF is the first method to integrate conditional diffusion models into particle filtering, enabling high-quality posterior sampling that produces more informative particles and significantly improves state estimation. The code is available at https://github.com/ZiyuNUS/DiffPF.
@@ I Introduction
Estimating the hidden state of a dynamic system from partial and noisy observations is a core challenge across a wide range of applications. Kalman filters and their extensions [thrun2005probabilistic] have long served as standard tools within the Bayesian filtering framework. However, their reliance on linearity and Gaussian noise assumptions severely limits their applicability to real-world systems that are often nonlinear, high-dimensional and exhibit multimodal behaviors.
@@
An alternative line of work, particle filters, avoids these restrictive assumptions by representing the posterior distribution with a set of weighted samples instead of relying on closed-form expressions. Bootstrap particle filter [gordon1993novel], one of the most commonly used algorithms, draws particles from the transition model and assigns weights based on observation likelihoods. Rao-Blackwellized particle filter [murphy2001rao] further improves efficiency by analytically marginalizing part of the latent state using parametric inference, thereby reducing sampling variance. However, these methods still rely on manually designed or limited-capacity proposals and are prone to sample degeneracy, particularly in high-dimensional or perceptually complex domains.
@@
Recent developments in deep learning have introduced data-driven alternatives to traditional state estimation methods. Deep state-space models (DSSMs) [rangapuram2018deep] provide a unified framework capable of capturing nonlinear latent dynamics and processing high-dimensional inputs, making them particularly suited for complex real-world environments. Differentiable Particle Filters (DPFs), a class of methods that embed sample-based filtering into DSSMs [chen2023overview, jonschkowski2018differentiable], enable end-to-end learning of recursive Bayesian inference over learned dynamics and observation models. However, most DPFs still inherit several limitations from classical particle filtering. The proposal distributions used in DPFs are often either hand-crafted or limited in capacity, making it difficult to match the true filtering posterior [karkus2018particle]. Furthermore, they rely on importance sampling, which necessitates explicit weighting of particles—a process that becomes increasingly unreliable in high-dimensional or highly multimodal settings due to particle degeneracy.
Figure 1: A conditional diffusion model parameterized by a U-Net is used to model the filtering posterior distribution and generate equally weighted particles samples via iterative denoising.
@@
To overcome these challenges, we employ diffusion models as powerful implicit samplers, capable of capturing rich and multimodal posterior distributions. This allows for a more expressive and learnable formulation of the state update step within the filtering process. Specifically, as illustrated in Fig. 1, DiffPF employs a diffusion model for state update, conditioned on both the predicted particles and the current observation to generate refined posterior samples. DiffPF departs from particle filtering by eliminating the reliance on importance sampling and explicit proposal distributions. Instead, it leverages a diffusion model to sample directly from the filtering posterior, enabling accurate inference without the need for particle weighting. This sampling-based yet weight-free approach naturally avoids particle degeneracy and, consequently, eliminates the need for the non-differentiable resampling step, while supporting multimodal and high-dimensional distributions.
Our main contributions are summarized as follows:
@@ II-A Particle filter
Particle filters (PFs) are a class of sequential Monte Carlo methods widely used for recursive state estimation in nonlinear and non-Gaussian systems [djuric2003particle]. Classical variants, such as the Bootstrap Filter [gordon1993novel], follow the Sampling-Importance-Resampling framework [godsill2019particle], which approximates the filtering distribution with weighted particles. While theoretically general, traditional PFs often suffer from particle degeneracy—where most weights collapse to near zero—and from the difficulty of designing effective proposal distributions, especially in high-dimensional or perceptually complex environments. Various extensions [godsill2019particle] have been developed to mitigate these issues by improving sampling efficiency, yet they typically remain limited by the same challenges of weight collapse and proposal design in complex state spaces.
@@ II-B Differentiable Particle Filters
The aforementioned traditional methods often assume known system dynamics or observation models, which often do not hold in real-world scenarios [chen2024normalizing]. Differentiable particle filters [chen2023overview, jonschkowski2018differentiable, karkus2018particle, corenflos2021differentiable, scibior2021differentiable, rosato2022efficient, kloss2021train, ma2020particle] address this by integrating neural networks into the filtering process, enabling end-to-end learning from data.
Most DPFs model transition dynamics by mapping previous states to the next latent state using either fully connected networks[jonschkowski2018differentiable, karkus2018particle], recurrent architectures such as LSTMs and GRUs[scibior2021differentiable, ma2020particle], or fixed parametric forms with learnable noise parameters[kloss2021train]. Observation models are typically implemented as neural networks that either output the parameters of a predefined distribution (e.g., the mean and variance of a Gaussian) [ma2020particle] or produce unnormalized scores interpreted as pseudo-likelihoods [corenflos2021differentiable]. While effective in simple scenarios, these designs are often restricted to specific distribution families and struggle to represent complex or multimodal posteriors, limiting the expressiveness of the resulting filtering distributions and motivating the need for a more flexible and general framework.
The proposal distribution, which guides particle sampling, is typically handcrafted [le2018auto, naesseth2018variational] or implemented as a neural network conditioned on the current observation and possibly past states. In many DPFs, the transition model is reused as the proposal, and particles are often generated deterministically via regression rather than sampling, which leads to particle degeneracy [bickel2008sharp], reduces particle quality, and constrains the representational capacity of the filtering distribution. Although [chen2024normalizing] proposes using normalizing flows to sample from the proposal distribution, the method remains confined to the importance sampling framework and suffers from accumulated estimation errors due to sampling from multiple distributions simultaneously. To enable gradient-based learning, DPFs must address the non-differentiability of traditional resampling operations. Several methods replace hard resampling with differentiable approximations [zhu2020towards], such as soft-resampling [karkus2018particle] and entropy-regularized optimal transport solvers [corenflos2021differentiable]. While these approximations allow gradients to flow through resampling steps, they often introduce additional computational complexity or require task-specific tuning, and particle degeneracy can still persist.
@@ II-C Diffusion Models for Probabilistic Modeling
Generative modeling has gained increasing attention across diverse domains [liu2025filtersallgeneralistfilter, lind2025normalizingflowscapablevisuomotor]. Recent advances in generative modeling have highlighted diffusion models as effective tools for capturing intricate, multimodal distributions across a range of data types. These models operate by progressively denoising latent variables [sohl2015deep, ho2020denoising], making them particularly adept at representing uncertainty and complex data structures.
In the domain of robotics, such approaches as Diffusion Policy [chi2023diffusion] and NoMaD [sridhar2024nomad] have demonstrated the effectiveness of conditional diffusion models in generating action sequences by capturing the inherently multi-modal nature of action distributions [wang2022diffusion]. Meanwhile, diffusion models have gained significant attention for their effectiveness in trajectory generation [carvalho2023motion, liang2023adaptdiffuser]. Building on the success of these tasks, we explore their potential in state estimation. Existing attempts include introducing diffusion models into Kalman filters [wan2025dnd]. We further develop a novel sample-based framework that integrates diffusion models into differentiable particle filtering, and propose end-to-end efficient training algorithms to facilitate the filter design.
Figure 2: At time $t$, the estimated particles from time t1t{-}1 are propagated through the process model to obtain a prior distribution over the current state. Simultaneously, the observation $\bm{o}_{t}$ is encoded into a feature representation. These two sources of information jointly condition a diffusion model, which iteratively refines a set of noisy latent samples to generate equally weighted particles that approximate the filtering posterior. Unlike DnD Filter [wan2025dnd], which uses a diffusion model in a Kalman filter–like manner to fuse prior and observation into a single trajectory, DiffPF maintains a full set of particles to explicitly represent the posterior distribution.
@@ III Method
In sequential inference, a Bayesian filter aims to recursively estimate the current state $\bm{x}_{t}$ via the posterior distribution $\text{post}(\bm{x}_{t}):=p(\bm{x}_{t}\mid\bm{o}_{1:t},\bm{a}_{1:t})$. Under the Markov assumption, all relevant information from past observations $\bm{o}_{1:t}$ and actions $\bm{a}_{1:t}$ is captured by the current posterior $\text{post}(\bm{x}_{t})$. By exploiting the conditional independence of observations and actions given the latent state, the inference proceeds recursively in two steps—prediction and update step [sarkka2023bayesian, jonschkowski2018differentiable]. As illustrated in Figure 2, our method follows this two-step Bayesian filtering procedure and represents the posterior $\text{post}(\bm{x}_{t})$ using a set of equally weighted particles $\bm{\chi}_{t}:=\{\bm{x}_{t}^{(i)}\}_{i=1}^{N}$. The remainder of this section details each component of the proposed framework.
III-A Prediction and Perception Modeling
Given a set of $N$ particles $\{\bm{x}_{t-1}^{(i)}\}_{i=1}^{N}$ representing the posterior at time t1t{-}1, we apply a process model to obtain predicted particles ${\hat{\bm{\chi}}_{t}}:=\{\hat{\bm{x}}_{t}^{(i)}\}_{i=1}^{N}$ for the current step. Specifically, each particle is propagated through a learnable or given process model $f_{\text{dyn}}$ as:
@@ III-A Prediction and Perception Modeling
where $\bm{a}_{t}$ is the action at time $t$. Through this step, the relevant information from the action sequence $\bm{a}_{1:t}$ and past observations $\bm{o}_{1:t-1}$ is encoded into the predicted particle set $\{\hat{\bm{x}}_{t}^{(i)}\}_{i=1}^{N}$. The process model $f_{\text{dyn}}$ can be defined using either parametric forms based on system knowledge, or learned from data using neural networks such as Multi-Layer Perceptrons (MLPs), stochastic models [jospin2022hands], or generative approaches like normalizing flows [papamakarios2021normalizing].
@@ III-B Update Step
Update step refines the prior belief over the current state by incorporating the latest observation. In conventional DPFs, the update is typically implemented by assigning importance weights to particles based on an observation model, followed by resampling to mitigate particle degeneracy. However, these procedures often require carefully designed likelihood functions and involve non-differentiable operations. In contrast, our method performs the update implicitly through a conditional diffusion model. Specifically, we make an assumption that
@@
This assumption is justified because $\bm{c}_{t}$ captures all temporal information from $\bm{o}_{1:t}$ and $\bm{a}_{1:t}$. Instead of sampling directly from the intractable posterior distribution $p(\bm{x}_{t}\mid\bm{o}_{1:t},\bm{a}_{1:t})$, we choose to sample from the approximated distribution $p(\bm{x}_{t}\mid\bm{c}_{t})$. This yields a set of particles $\{\bm{x}_{t}^{(i)}\}_{i=1}^{N}$, from which we can obtain:
$\text{post}(\bm{x}_{t})$ $\approx\frac{1}{N}\sum_{i=1}^{N}\delta(\bm{x}_{t}-\bm{x}_{t}^{(i)}),$ (5)
where $\delta(\cdot)$ denotes the Dirac function. This indicates that all sampled particles share equal weights, thereby eliminating the need for weighting or resampling.
Diffusion models have been shown to approximate arbitrarily complex probability distributions by learning time-dependent score functions that guide the reversal of a noise-perturbed process[song2020denoising]. In our case, a conditional diffusion model is trained to approximate $p(\bm{x}_{t}\mid\bm{c}_{t})$ and generate particle samples. Specifically, we instantiate a reverse diffusion process with a Denoising Diffusion Probabilistic Model (DDPM), which iteratively refines a random Gaussian sample toward the target posterior distribution. We begin by drawing an initial Gaussian sample:
@@
where $K$ is the total number of reverse diffusion steps. At each step $k$, the model predicts the noise component $\bm{\epsilon}_{\theta}(\bm{x}_{t,k}^{(i)},k,\bm{c}_{t})$, which subsequently serves to infer a cleaner approximation of the sample at timestep k1k-1:
@@
$=\beta(\bm{x}_{t,k}^{(i)}+\gamma\epsilon_{\theta}(\bm{x}_{t,k}^{(i)},k,\bm{c}_{t}))+\sigma\bm{z}.$
Here $\bm{z}\sim\mathcal{N}(0,I)$, and $\bm{x}_{t,k}^{(i)}$ refers to the $i$-th corrupted sample at diffusion step $k$. $\alpha$, $\bar{\alpha}$, $\sigma$} are fixed noise schedule parameters as defined in [chan2024tutorial], while $\beta$ and $\gamma$ are scaling terms introduced for simplicity. This stochastic refinement continues recursively until $k=0$, yielding the final particle sample $\bm{x}_{t,0}^{(i)}=\bm{x}_{t}^{(i)}.$ This reverse diffusion process is repeated $N$ times to produce the final particle set $\bm{\chi}_{t}=\{\bm{x}_{t}^{(i)}\}_{i=1}^{N}$. As all particles are equally weighted, the final estimate is given by the average over the particle set
$\bm{x}_{t}=\frac{1}{N}\sum_{i=1}^{N}\bm{x}_{t}^{(i)}.$ (8)